hw is the cureent in the solenoid created when the magnetic lines of force linked with it change with respect to time?????????????/

Grade:

1 Answers

vikas askiitian expert
13 years ago

if the rate of change of magnetic flux lincked with a coil is dB/dt is given then

from the faradays second law

                         d@/dt = dB/dt (A)                      (A is area of cross section and @ i is flux )

                          e=AdB/dt

                         I=(AdB/dt)/R                   (R is the resistance of coil)